# DermaCon-IN: A Multi-concept Annotated Dermatological Image Dataset of Indian Skin Disorders for Clinical AI Research

### v1

@@ 6.1 Standard architectures
DermaCon-IN comprises high-resolution clinical photographs with multiple co-occurring lesions, varied anatomical regions, and hierarchically structured multi-label annotations. We selected architectures based on their complementary modelling strengths to benchmark model performance under these conditions. Convolutional neural networks such as ResNet50 [13], DenseNet121 [16], and EfficientNet (EffNet-B4_b4[35] are effective in capturing localized texture patterns and edge-level features, owing to their convolutional inductive biases and limited receptive fields. To complement this, we incorporated Vision Transformer (ViT) architectures [6], which leverage self-attention to relate spatially distant regions within an image. The ViT variants evaluated includes ViT-Base (ViT-B/16-224 [6], ViT-B/16-384[6], MaxViT-Base (MaxViT-B/512[38], and Swin-Base (Swin-B/4W12-384[24]. Among these, the Swin Transformer consistently achieved the best results for Main class prediction across evaluation metrics, demonstrating improved handling of both multi-class classification and class imbalance. Table 2 summarizes the performance of all models considered. Swin Transformer’s shifted window mechanism enables efficient modeling of non-contiguous regions, while its hierarchical representation captures both fine-grained lesion details and broader spatial patterns. These traits align closely with our dataset. We believe this alignment contributed to Swin’s better performance.
Based on these observations, we adapted the Swin-B/4W12-384 [24] variant of the Swin Transformer as the backbone for subsequent analysis and in concept bottleneck models (CBMs) [22] as shown in Table 3. The model was initialized with weights pretrained on ImageNet-22k and fine-tuned end-to-end on our dataset. Input images were resized and padded to $512\times 512$ for further classification. We performed a stratified, subject-wise 80:20  split over Sub Class and reported all the results with the same split. We adapted weighted sampling strategies to handle class imbalance, and details of which are discussed in Supplementary Sec. B.
@@ 6.2 Concept Bottleneck Modeling for Interpretablity
Given an input image $x$, the Swin-Transformer encoder $E_{\theta}$ yields a latent representation $z=E_{\theta}(x)$. A linear projection maps $z$ to concept logits $c^{\ell}\in\mathbb{R}^{B+D}$, which are then passed through a sigmoid activation to obtain the interpretable concept vector $c=[c^{\text{bp}},\,c^{\text{ld}}]=\sigma(c^{\ell})\in[0,1]^{B+D},$ where $c^{\text{bp}}$ denotes $B$ body-part concepts and $c^{\text{ld}}$ denotes $D$ lesion-descriptor concepts. While both $c$ and $c^{\ell}$ is used for interpretability and concept supervision, the downstream classifier $f$ operates on the concept logits $c^{\ell}$ to produce task logits $y=f(c^{\ell})$, predicting either a Main-Class (MC) or Sub-Class (SC) label. This architecture is presented in Figure 7(A).
@@
To quantify the contribution of each concept group we trained two ablated models: CBM-D, which keeps only lesion-descriptor concepts $\smash{c^{\text{ld}}}$, and CBM-B, which keeps only body-part concepts $\smash{c^{\text{bp}}}$, both predicting MC labels. Each single–stream variant retained high accuracy (Table. 2) demonstrating that the two concept families are independently learnable.
By contrast, the full CBM represents the true clinical diagnostic fidelity, in which both $\smash{c^{\text{bp}}}$ and $\smash{c^{\text{ld}}}$ co-exist in the bottleneck, achieved performance comparable to the individual concept streams, but revealed a systematic imbalance in activation: In many samples, only one concept group (typically descriptors) fired strongly, whereas the other (body parts) was under-activated (Fig. 8, bottom row). This led to a modest but consistent drop in overall accuracy, pointing to a representational bottleneck whereby competition for limited capacity biases the model toward a single semantic stream. These observations emphasise the need for improved multi-concept learning mechanisms that can balance several concept families simultaneously.
@@
We explore two designs for joint SC–MC prediction (Figure. 7(B)):
@@
    Type 1 — cascade. Concepts first predict sub-classes via $y1=f(c)$; these logits are then mapped to main classes through a second head $y2=g(y1)$, ensuring taxonomy consistency by construction.
@@
To further assess whether the model’s predictions are semantically grounded, we evaluate the alignment between its learned class-concept weights (MC branch of Type-2) and the statistical relevance (Pearson correlation, Section 4) of each concept to the class labels in the dataset (Figure 10). Alignment varies notably across diagnostic categories. Pigmentary Disorders and Keratinisation Disorders show strong Spearman correlations and statistically significant p-values ($p<0.05$), suggesting that the model reliably prioritizes clinically meaningful features for these classes. In contrast, Neoplasms and Tumors show weak or negative alignment, indicating reliance on non-semantic or latent cues, possibly due to low representation in the dataset. Whereas, No Definite Diagnosis shows negative correlations as desired. Classes like Skin Appendageal Disorders, etc. exhibit partial alignment.
@@
DermaCon-IN captures real-world dermatological presentations from the South Indian population, offering concept-level annotations such as body parts and disease descriptors. It addresses gaps in skin tone diversity and regional disease patterns, serving both as a region-specific benchmark and a valuable addition to global dermatological datasets. By enabling evaluation beyond labels, it supports clinically grounded, interpretable learning—a step closer towards clinical deployment.
@@
  • Curriculum and re-sampling strategies: oversampling under-represented concept combinations (e.g. body-part signals within neoplasm classes) to equalise learning pressure across concept space.
Broader integration. Because DermaCon-IN offers strong coverage of South-Indian skin tones and disease spectra, its concept annotations complement existing global datasets. Merging these resources will enable training of larger, more broadly representative models—paving the way toward a unified foundation-level representation for dermatological imaging.
@@
The experiments were run on 4 GPUs of Nvidia-A6000, each of $42$GB RAM. The dataset sizes around $\sim\!4$ GB. The code used in this work is available at GitHub.
@@
This work was not supported by any external funding. All contributors were involved out of self-motivation and shared interest in advancing dermatological AI research. We would also like to thank Dr. Deepanshu Bansal for his contributions to the dataset labelling process.
@@ A.2 Descriptor Design and Clinical Alignment
We annotate $5{,}450$ dermatological images using $47$ carefully selected descriptors encompassing primary lesions, secondary changes, pigmentary alterations, vascular anomalies, and surface/nail abnormalities. The taxonomy is rooted in established dermatological frameworks [3], ensuring clinical interpretability and consistency with medical curricula and diagnostic practice. While many descriptors overlap with those used in SkinCon [5] (e.g., papule, plaque, vesicle), our set incorporates necessary refinements based on regional prevalence and diagnostic relevance. Descriptors with high clinical utility but low frequency (e.g., poikiloderma, gray, salmon) were retained to preserve diagnostic granularity. The descriptors such as translucent, friable, and dome-shaped lacked distinct visual features in $2D$ clinical images and so were not included in the list of descriptors, which were otherwise part of SkinCon descriptors.
@@
Table 4: Complete Concept Descriptor Table with Definitions and Reasoning.
S.No Concept Explanation Reason for Choosing the Descriptor
@@
6 Burrow It is a thin, wavy, slightly raised line on the skin, typically found between the fingers or toes, caused by the scabies mite burrowing under the surface. Highly relevant in scabies, which has significant endemic prevalence in India.
@@
11 Discolored Nail Alteration in nail color due to trauma, infection, pigmentation, or systemic disease. Indicator of systemic illness, fungal infections, or trauma-related disorders.
@@
45 White (Hypopigmentation) Lighter than normal skin color due to loss or reduction of melanin. Central to vitiligo, pityriasis alba, and tinea versicolor which are frequent in India.

### v2

@@ 6.1 Standard architectures
DermaCon-IN comprises high-resolution clinical photographs with multiple co-occurring lesions, varied anatomical regions, and hierarchically structured multi-label annotations. We selected architectures based on their complementary modelling strengths to benchmark model performance under these conditions. Convolutional neural networks such as ResNet50 [13], DenseNet121 [17], and EfficientNet (EffNet-B4[38] are effective in capturing localised texture patterns and edge-level features, owing to their convolutional inductive biases and limited receptive fields. To complement this, we incorporated Vision Transformer (ViT) architectures [6], which leverage self-attention to relate spatially distant regions within an image. The ViT variants evaluated includes ViT-Base (ViT-B/16-224 [6], ViT-B/16-384[6], MaxViT-Base (MaxViT-B/512[41], and Swin-Base (Swin-B/4W12-384[25]. Among these, the Swin Transformer consistently achieved the best results for Main class prediction across evaluation metrics, demonstrating improved handling of both multi-class classification and class imbalance. Table 2 summarises the performance of all models considered. Swin Transformer’s shifted window mechanism enables efficient modeling of non-contiguous regions, while its hierarchical representation captures both fine-grained lesion details and broader spatial patterns. These traits align closely with our dataset. We believe this alignment contributed to Swin’s better performance.
Based on these observations, we adapted the Swin-B/4W12-384 [25] variant of the Swin Transformer as the backbone for subsequent analysis and in concept bottleneck models (CBMs) [23] as shown in Table 3. The model was initialised with weights pretrained on ImageNet-22k and fine-tuned end-to-end on our dataset. Input images were resized and padded to $512\times 512$ for further classification. We performed a stratified, subject-wise 80:20  split over Sub Class and reported all the results with the same split. We adapted weighted sampling strategies to handle class imbalance, and details of which are discussed in Supplementary Sec. B.
@@ 6.2 Concept Bottleneck Modeling for Interpretablity
Given an input image $x$, the Swin-Transformer encoder $E_{\theta}$ yields a latent representation $z=E_{\theta}(x)$. A linear projection maps $z$ to concept logits $c^{\ell}\in\mathbb{R}^{B+D}$, which are then passed through a sigmoid activation to obtain the interpretable concept vector $c=[c^{\text{bp}},\,c^{\text{ld}}]=\sigma(c^{\ell})\in[0,1]^{B+D},$ where $c^{\text{bp}}$ denotes $B$ body-part concepts and $c^{\text{ld}}$ denotes $D$ lesion-descriptor concepts. While both $c$ and $c^{\ell}$ are used for interpretability and concept supervision, the downstream classifier $f$ operates on the concept logits $c^{\ell}$ to produce task logits $y=f(c^{\ell})$, predicting either a Main-Class (MC) or Sub-Class (SC) label. This architecture is presented in Figure 7(A).
@@
To quantify the contribution of each concept group, we trained two ablated models: CBM-D, which keeps only lesion-descriptor concepts $\smash{c^{\text{ld}}}$, and CBM-B, which keeps only body-part concepts $\smash{c^{\text{bp}}}$, both predicting MC labels. Each single–stream variant retained high accuracy (Table 2), demonstrating that the two concept families are independently learnable.
By contrast, the full CBM represents the true clinical diagnostic fidelity, in which both $\smash{c^{\text{bp}}}$ and $\smash{c^{\text{ld}}}$ co-exist in the bottleneck, achieved performance comparable to the individual concept streams, but revealed a systematic imbalance in activation: In many samples, only one concept group (typically descriptors) fired strongly, whereas the other (body parts) was under-activated (Fig. 8, bottom row). This led to a modest but consistent drop in overall accuracy, pointing to a representational bottleneck whereby competition for limited capacity biases the model toward a single semantic stream. These observations emphasize the need for improved multi-concept learning mechanisms that can balance several concept families simultaneously.
@@
We explore two designs for joint SC–MC prediction (Figure 7(B)):
@@
    Type 1 — cascade. Concepts first predict sub-classes via $y1=f(c)$. These logits are then mapped to the main classes through a second head $y2=g(y1)$, ensuring taxonomy consistency by construction.
@@
To further assess whether the model’s predictions are semantically grounded, we evaluate the alignment between its learned class-concept weights (MC branch of Type-2) and the statistical relevance (Pearson correlation, Section 4) of each concept to the class labels in the dataset (Figure 10). Alignment varies notably across diagnostic categories. Pigmentary Disorders and Keratinisation Disorders show strong Spearman correlations and statistically significant p-values ($p<0.05$), suggesting that the model reliably prioritises clinically meaningful features for these classes. In contrast, Neoplasms and Tumors show weak or negative alignment, indicating reliance on non-semantic or latent cues, possibly due to low representation in the dataset. Whereas No Definite Diagnosis shows negative correlations as desired. Classes like Skin Appendageal Disorders, etc. exhibit partial alignment.
@@
DermaCon-IN captures real-world dermatological presentations from the South Indian population, offering concept-level annotations such as body parts and disease descriptors. It addresses gaps in skin tone diversity and regional disease patterns, serving both as a region-specific benchmark and a valuable addition to global dermatological datasets. By enabling evaluation beyond labels, it supports clinically grounded, interpretable learning, a step closer towards clinical deployment.
@@
  • Curriculum and re-sampling strategies: oversampling under-represented concept combinations (e.g. body-part signals within neoplasm classes) to equalize learning pressure across concept space.
Broader integration. Because DermaCon-IN offers strong coverage of South-Indian skin tones and disease spectra, its concept annotations complement existing global datasets. Merging these resources will enable training of larger, more broadly representative models, paving the way toward a unified foundation-level representation for dermatological imaging. Looking ahead, we will expand DermaCon-IN as versioned releases sourced from additional centers across India, thereby deepening geographic, phenotypic, and skin-tone coverage while preserving curation standards. In parallel, we plan to add pixel-level lesion masks for a subset of images, enabling segmentation and tighter alignment between concepts and spatial evidence. We hope this resource will move dermatology AI closer to responsible clinical use.
@@
The experiments were run on 4 GPUs of Nvidia-A6000, each of 4242GB RAM. The dataset sizes around $\sim\!4$ GB. The code used in this work is available at GitHub.
@@
This work was not supported by any external funding. All contributors were involved out of self-motivation and shared interest in advancing dermatological AI research.
@@ A.2 Descriptor Design and Clinical Alignment
We annotate 5,4505{,}450 dermatological images using $47$ carefully selected descriptors encompassing primary lesions, secondary changes, pigmentary alterations, vascular anomalies, and surface/nail abnormalities. The taxonomy is rooted in established dermatological frameworks [3], ensuring clinical interpretability and consistency with medical curricula and diagnostic practice. While many descriptors overlap with those used in SkinCon [5] (e.g., papule, plaque, vesicle), our set incorporates necessary refinements based on regional prevalence and diagnostic relevance. Descriptors with high clinical utility but low frequency (e.g., poikiloderma, gray, salmon) were retained to preserve diagnostic granularity. The descriptors such as translucent, friable, and dome-shaped lacked distinct visual features in $2D$ clinical images and so were not included in the list of descriptors, which were otherwise part of SkinCon descriptors.
@@
Table 5: *
@@
6 Burrow It is a thin, wavy, slightly raised line on the skin, typically found between the fingers or toes, caused by the scabies mite burrowing under the surface. Highly relevant in scabies, which has a significant endemic prevalence in India.
@@
11 Discolored Nail Alteration in nail color due to trauma, infection, pigmentation, or systemic disease. an Indicator of systemic illness, fungal infections, or trauma-related disorders.
@@
45 White (Hypopigmentation) Lighter than normal skin color due to loss or reduction of melanin. Central to vitiligo, pityriasis alba, and tinea versicolor, which are frequent in India.